OpenStudy (meilendurcer):

The science club sold 25 T-shirts last week and 18 T-shirts this week. What is the percentage of decrease in the amount of T-shirts sold? Answers; 72% 28% 39% 36%

OpenStudy (anonymous):

A decrease of about 16% 25+18= 43 25/43= 58% 18/43=41% 58-41= 16%

OpenStudy (anonymous):